# Conformal Alignment: Knowing When to Trust Foundation Models with Guarantees

### v1

@@ C.2 CXR report generation
Once fine-tuned, the model is treated as fixed for generating reports. The data not used in fine-tuning are used as the training, calibration, and test data in our experiments. For report generation, we use the default configurations (max_length=512, do_sample=True) and the input of the vision-language model consists of pixel values of images (resized to $224\times 224$ pixels) and tokenized prompts (tokenized via the same GPT2 model). For each image, we generate $M=10$ reports independently, with which we treat the first output as $f(X_{i})$ and utilize the $M$ outputs to calculate the features (similary-based scores) for training the alignment predictor. To determine if a generated report is admissible, we follow the practice in [33] and use the CheXbert model [41] to convert both the generated and reference reports into $14$-dimensional vectors. For each coordinate, the entry falls into the categories “Blank”, “Positive”, “Negative” or “Uncertain”, with which we further convert them to binary vectors with each entry indicating “Positive” or not. The alignment score $A_{i}\in\{0,1\}$ is defined to be the indicator of whether there are at least $12$ matched coordinates between the two binary vectors. Similar to the QA task, we modify the pipeline in [27] to calculate the uncertainty and confidence scores, in which we utilize an off-the-shelf DeBERTa-large model [11] as the NLI classifier to calculate similarities.
@@ D.1 FDR/Power for TriviaQA dataset with alternative classifiers
We present the FDR/power plots when applying Conformal Alignment to the TriviaQA dataset and the alignment predictor is trained with random forests (Figure 8) and XGBoost (Figure 9), fixing $\gamma_{1}=0.2$ and $\gamma_{2}=0.5$. These figures are similar to Figure 3 (which uses logistic regression to train the alignment predictor) in the main text. We observe similar messages as those in Section 4: the FDR is always tightly controlled with satisfactory power. Regarding the choice of $|\mathcal{D}|$, we similarly see that $500$ high-quality samples suffice for powerful deployment of our method.
@@ D.2 FDR/Power on CoQA dataset
We present the FDR/power plots for the CoQA dataset when the alignment predictor is trained with logistic regression (Figure 10), random forest (Figure 11), and XGBoost (Figure 12), fixing $\gamma_{1}=0.2$ and $\gamma_{2}=0.5$. We again observe similar messages as those in Section 4: the FDR remains tightly controlled, and a sample size of $|\mathcal{D}|=500$ suffices for powerful deployment of our method.

### v3

@@ B.1 Comparison with selective prediction
B.2 Comparison with conformalized selection
Our work instantiates conformalized selection [18] for the reliable deployment of foundation model outputs. We detail our contributions as follows. First, we justify that selecting reliable outputs with FDR control is a practically reasonable criterion for downstream tasks. Second, we formulate the alignment problem as a selective prediction problem that can be tackled with conformal selection, adapting conformal selection to the current setting. Note that while conformal selection is an established method, its current application in job hiring and drug discovery requires numerical outputs and predictions; the task of aligning foundation model outputs, however, does not come with a clear numerical output. To make conformal selection applicable to our purpose, we formulate the alignment status as a numerical indicator that could be inferred with reference information, and then decide how to train a prediction model for the alignment score, which can therefore be used in conjunction with the conformal selection framework. Our third contribution is the implementation of the pipeline. In this regard, we conduct extensive numerical experiments, finding that using a lightweight prediction model such as logistic regression or random forests with popular heuristic (uncalibrated) uncertainty measures is often sufficient for identifying reliable outputs. This provides a concrete and easy-to-use pipeline for practical use.
@@ C.2 CXR report generation
Once fine-tuned, the model is treated as fixed for generating reports. The data not used in fine-tuning are used as the training, calibration, and test data in our experiments. For report generation, we use the default configurations (max_length=512, do_sample=True) and the input of the vision-language model consists of pixel values of images (resized to $224\times 224$ pixels) and tokenized prompts (tokenized via the same GPT2 model). For each image, we generate $M=10$ reports independently, with which we treat the first output as $f(X_{i})$ and utilize the $M$ outputs to calculate the features (similarity-based scores) for training the alignment predictor. To determine if a generated report is admissible, we follow the practice in [36] and use the CheXbert model [44] to convert both the generated and reference reports into $14$-dimensional vectors. For each coordinate, the entry falls into the categories “Blank”, “Positive”, “Negative” or “Uncertain”, with which we further convert them to binary vectors with each entry indicating “Positive” or not. The alignment score $A_{i}\in\{0,1\}$ is defined to be the indicator of whether there are at least $12$ matched coordinates between the two binary vectors. Similar to the QA task, we modify the pipeline in [29] to calculate the uncertainty and confidence scores, in which we utilize an off-the-shelf DeBERTa-large model [13] as the NLI classifier to calculate similarities.
@@ D.1 FDR/Power for TriviaQA dataset with alternative classifiers
We present the FDR/power plots when applying Conformal Alignment to the TriviaQA dataset and the alignment predictor is trained with random forests (Figure 8) and XGBoost (Figure 9), fixing $\gamma_{1}=0.2$ and $\gamma_{2}=0.5$. These figures are similar to Figure 3 (which uses logistic regression to train the alignment predictor) in the main text. We observe similar messages as those in Section 4: the FDR is always tightly controlled with satisfactory power. Regarding the choice of $|\mathcal{D}|$, we similarly see that $500$ high-quality samples suffice for the powerful deployment of our method.
@@ D.2 FDR/Power on CoQA dataset
We present the FDR/power plots for the CoQA dataset when the alignment predictor is trained with logistic regression (Figure 10), random forest (Figure 11), and XGBoost (Figure 12), fixing $\gamma_{1}=0.2$ and $\gamma_{2}=0.5$. We again observe similar messages as those in Section 4: the FDR remains tightly controlled, and a sample size of $|\mathcal{D}|=500$ suffices for the powerful deployment of our method.
@@ D.5 Real examples in QA
D.6 Empirical evaluation of feature importance
In Figure 5 and 7, we quantify the importance of individual features in the alignment predictor via the ROC curve. In this section, we use the Shapley value as well as the model-based score to evaluate the individual feature importance more directly.
We used the XGBoost functionality to compute the importance scores in Figure 19(d). We also computed the Shapley value of individual features to measure their importance based on the prediction models we use in the manuscript. As the Shapley values are model-agnostic, we present results for three choices of $g(X)$ with different base classifiers, including logistic regression, random forests, and XGBoostRF, in Figure 19(a), 19(b), and 19(c). Although the exact order of features varies in four plots, there are features that have consistent dominating effects in all figures, e.g. Deg(J) and EigV(J) based on the Jaccard similarity. In addition, NumSets as the feature in alignment score predictors tends to exhibit an effect close to zero in all four figures. We should note that the aforementioned findings are also revealed in Figure 5 in the paper, where predicted scores with EigV(J), Deg(J), and Ecc(J) have higher power and that with NumSets is much less powerful. We thank the reviewer again for raising this question for more comparison of feature importance in the alignment score predictor.
(a) Shapley Values: logistic regression.
(b) Shapley Values: Random Forests
(c) Shapley Values: XGBRF
(d) Feature importance for XGBRF extracted from sklearn.
Figure 19: Measures of Feature importance (Dataset: TriviaQA, LLM: LLaMA-2-13B-chat).
D.7 Comments on FDR guarantee and comparison to other frameworks
The major distinction between our method and other applications of conformal prediction on the foundation models is the type of guarantee we pursue. This is related to the following question: how can we use the outputs of these algorithms in practice? In short, (a) For [36], it is unclear how to pick one answer from the multiple answers in a way that does not hinder validity. (b) Conformal actuality [32] makes the original outputs less specific, which can hinder downstream use.
To demonstrate (a), we conducted additional experiments for more clear comparison. With the TriviaQA dataset and the base model LLaMA-2-13B-chat, we follow the procedure in [36] and have the following examples:
  • Example 1: Question: In tennis, losing two sets 6-0 is known as a double what? True answer: Bagel. Generated set: {‘bagel’}.
  • Example 2: Question: What was the name of the brothel in The Best Little Whorehouse in Texas? True answer: Chicken ranch. Generated set: {‘miss monas’, ‘chicken ranch’, ‘miss mona’s’, ‘miss monas bordello’}.
As we can see in Example 1, the generated set is very informative and can be used directly. However, although the generated set in the second example has a high probability of covering the true answer, it contains very different answers and is potentially confusing in practice. In other words, whether “Chicken ranch” or “Miss Monas” is correct is unclear to the user.
In addition, as is shown in Figure 20 (left panel), we present the proportion that the size of generated size is $k$ for different values of $k\in\mathbb{N}$, which shows that there are more than $40\%$ of sets having no less than $4$ distinct items. In addition, Figure 20 (right panel) shows that in nearly $30\%$ of the questions, the admissible answers (with $\texttt{rougeL}\geq 0.3$) take up less than $20\%$ of the uncertain set, which poses a challenge in the direct deployment of answers in the uncertain set.
Figure 20: Additional histograms for Conformal Language Modeling: sizes of uncertainty sets and proportions of admissible answers, where $C_{i}$ denotes the uncertainty set for question No.$i$ (Dataset: TriviaQA, LLM: LLaMA-2-13B-chat).